OpenStudy (anonymous):

I really need help guys. I will medal. Question: Gina has to spend 3/10 of her weekly income on food and board and 1/5 on travel to and from university and 1/8 on text books. What fraction of her income does she spend? b. Use fraction operations to show that 3/8 of her income remains at the end of each week? c. If her total weekly income is $520, calculate the amount that remains each week. D. Gina wants to save money for a holiday to bali at the end of the year. she estimates that if would cost her $2200 but she only has 14 weeks of work till then. Will she have enough money?

OpenStudy (anonymous):

I got part A... but do not understand part b,c,d

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

its easy the other way if u like?

OpenStudy (anonymous):

by %age method

OpenStudy (anonymous):

she spends 325$ weekly and saves 195$ and u can calculate further

OpenStudy (anonymous):

I'm still confused... @leon549

OpenStudy (anonymous):

so how do u calculate part "C" ?

OpenStudy (anonymous):

someone help meeeeeeeeeeee

OpenStudy (alekos):

(1-3/10-1/5-1/8)520 would be the answer to part c

OpenStudy (anonymous):

Thanks! What about D? @alekos

OpenStudy (anonymous):

For B, all the question is asking to show the steps to finding how much of Gina’s income is left over each week. Let g represent Gina’s income: If we subtract the expenses from her income, then we will get the ⅜ of her income that remains each week: \[g - \left( \frac{ 3g }{ 10 } + \frac{ 1g }{ 5 } + \frac{ 1g }{ 8 }\right)\] Then find the Least Common Multiple of these fractions multiply each fraction until all the denominators match, then add. \[g - \frac{ 5g }{ 8 }\] = \[\frac{ 3g }{ 8 }\] And that would be your answer (the whole thing though, because the question is asking for the process) For C, all it really wants you to do is replace “g” with $520 in the last part of B: \[\frac{ 3*520 }{ 8 }\] Then simplify Then For D, it wants to know if Gina's can save $2200 in 14 weeks, using the remainder of income. So what you would do is you would take the answer from part C and multiply it by 14 weeks, then compare and see if it's less than or greater than $2200. Hope that helps!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Although I see some question marks in your answer, can you fix that?

OpenStudy (anonymous):

I saw question marks, too, but after a while they went away. Try reloading the page.

OpenStudy (anonymous):

Ohhhh thanks! it's fixed! Thank you so much for helping me out. Can I ask you one more question?

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Gerald is training for the next Tour de France. He practices by doing a lap around a set route in his neighborhood. Each day his time is 5 minutes faster than the previous day. In total, he spent 10 hours in training over 4 days. Determine his lap time on the last day of practice. Hint: There 60 minutes in one hour.

OpenStudy (anonymous):

I have no idea on this....

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

Yes, I'm working on the problem :)

OpenStudy (anonymous):

THAAAAAAAAAAAAAAAAAANK UUUUUUUUUU <3

OpenStudy (anonymous):

Here's a hint: since he's 5 minutes faster every day, that also means that each day's practice time is 5 minutes shorter (because he does one lap a day).

OpenStudy (anonymous):

Using variables, this would mean: let p represent Gerald's practice time on the last day (the shortest time): The last day's practice time = p The 3rd day's practice time = p +5 The 2nd day's practice time = p +5 +5 The 1st day's practice time = p + 5 + 5 +5

OpenStudy (anonymous):

so just adding them all?

OpenStudy (anonymous):

hmmm.. never mind nope im wrong

OpenStudy (anonymous):

Well, actually, you're right. If you add all of those together, you will get the total practice time, which is 10 hours (according to the question). But since everything else is in minutes, you might want to multiply the 10 hours by 60 minutes first. :)

OpenStudy (anonymous):

I get 600 minutes?

OpenStudy (anonymous):

so ... P+P+5+P+5+5+P+5+5+5 = 600 ?

OpenStudy (anonymous):

Yes! Exactly!

OpenStudy (anonymous):

4P+ 30 = 600?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

Now 4P = 600-30 4P = 570 P = 570/4 P = 142.5 = his lap time?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Thaaaaaaaaaaaaaaaaank you so very much <33 you're such a big help... I just have one final question then I'll be done... so it says using numbers 120 and 75 answer the following a. Write each number as product of their prime factors. give your answer in index notation. I already got part B and C ... just confused on this one.

OpenStudy (anonymous):

You're very welcome, but you might want to post that one in another question because I'm eating dinner now, and it's going to take a long time. :)

OpenStudy (anonymous):

Oh thank you anyways! Really appreciate it! it's morning here in germany!

OpenStudy (anonymous):

Take Care!